# Millen Kanabar and Michael Gastpar School of Computer and Communication Sciences, EPFL

### v3

@@ IV-C1 An example with matching bounds
$=$ $\sum_{i=1}^{t}n_{i}\alpha_{i}\sup_{p}\max_{j}\bar{\mathbf{P}}_{i-1,({p})}\left\{\hat{P}_{i-1}[j]<p[j]-\nicefrac{{1}}{{4k}}\right\}$
$\leq$ $\sum_{i=1}^{t}a(i+1)^{\beta}\cdot\frac{\text{var}\left(\hat{P}_{i-1}\right)}{\left(\nicefrac{{1}}{{4k}}\right)^{2}}\leq\sum_{i=1}^{t}a(i+1)^{\beta}\cdot 16k^{2}\cdot r_{i-1,k}^{\ell_{2}^{2}}$
$\leq$ $\sum_{i=1}^{t}a(i+1)^{\beta}\cdot 16k^{2}\cdot\frac{\left(1-\frac{1}{k}\right)(\beta-\gamma+1)}{a\cdot i^{\beta-\gamma+1}}\leq 2^{\beta}\cdot 16k^{2}\cdot(\beta-\gamma+1)\sum_{i=1}^{t}i^{\gamma-1}$ (8)
@@
This inductively shows that the upper and lower bounds match in terms of order w.r.t. the stage $t$: for each stage $t$, the upper bound is within a constant factor of the lower bound, provided that order-optimal estimators are chosen for stages 00 through $t-1$. Moreover, there is a polynomially growing gap between the base and oracle-assisted minimax losses:
@@ IV-C2 An example where the upper and lower bound do not match
We should also note that stronger upper bounds, for example, of the form $h_{t}(\nicefrac{{1}}{{4k}})\leq\exp(-\lambda_{0}t)$, can be assumed by restricting the class of estimators appropriately; this would then make the estimators in Theorem 2 order optimal for every choice of parameters that satisfies the hypothesis condition. Bounds using the Chebyshev inequality remain the tightest without such assumptions.
@@ IV-D General regimes for data accumulation
completing the first part of the proof. If $\nicefrac{{1}}{{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}}=o(\alpha_{t})$, $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$ also diverges (since $\alpha_{t}\leq 1$), and one can find $t_{1}$ such that $g_{t}(\nicefrac{{1}}{{4k}})$ is smaller than $\epsilon$—independent of $k$—for $t\geq t_{1}$. The denominator of (12) can then be split into $(1+16k^{2}m)\sum_{i=0}^{t_{1}-1}n_{i}\alpha_{i}^{2}$ and $(1+\epsilon)\sum_{i=t_{1}}^{t}n_{i}\alpha_{i}^{2}$, the latter of which will dominate the former as $t$ grows large. ∎
@@
for each $i$ in $[0:t-1]$, for some $M>0$. Then there exists a constant $m>0$ such that the minimax loss at stage $t$ is bounded below as
@@
when the aforementioned estimators are selected for stages 00 to $t-1$.
The proof of this Proposition is deferred to Appendix D. The third condition here is satisfied by all sequences for which $n_{t}\alpha_{t}$ grows exponentially or slower (as can be seen by the Discrete Gronwall inequality [21]). This proposition solidifies the intuition on the lower bound when the real sampling probabilities $\alpha_{t}$ decay too quickly: the effective sample size is proportional to the square root of the sample size when in the oracle-assisted setting.
@@ IV-E Data replacement
In this section, we demonstrate the application of Corollary 4 in the data replacement setting where, at every stage $t$, the estimator only has access to samples generated in the current stage $X_{t}^{n}$. Note that $\alpha$ and $n$ do not change with $t$ in this setting, except at stage 00 where all $n$ samples are real. This setting has been explored in depth in the literature e.g. [6, 9, 7, 5, 11, 10, 14, 8].
@@
When $k$ is fixed, we find that since $M\leq\sqrt{2}$ trivially, $m\geq\nicefrac{{c}}{{n\alpha}}$ for some $c\geq 0$. This is especially stark in contrast to the oracle-assisted infimum, which, with probability $1$, is proportional to $\nicefrac{{1}}{{n}}$, corresponding to the case where all samples are real. Thus, in the data replacement setting, the ‘best-case’ minimax expected loss is of the same order as the ‘average’ oracle-assisted loss.
@@ Appendix A Proofs of the minimax lower bounds
We assume that the alphabet size $k$ is even. The worst-case loss for odd $k$ is lower bounded by the worst-case loss for $k-1$, since the supremum is over a smaller set of distributions.
The lower bound argument uses Assouad’s method with the following construction, widely known in the distribution estimation literature e.g. as described in [22, 23, 24, 25]. Consider the set of vectors $\mathcal{V}=\{+1,-1\}^{k/2}$. With each vector $v\in\mathcal{V}$, associate the distribution
@@
The following Lemma will be useful in the proof of the Theorem:
@@
We will now prove Theorem 1 by deriving an upper bound on the KL divergences between the distributions $\bar{\mathbf{P}}_{t,(p_{v})}$ and $\bar{\mathbf{P}}_{t,(p_{v-2e_{j}})}$ for each $v\in\mathcal{V}$.
@@
Using the chain rule for KL divergences, whenever $j\leq\nicefrac{{k}}{{2}},v[j]=+1$,
$D\Big{(}\bar{\mathbf{P}}_{t,(p_{v})}\Big{\|}\bar{\mathbf{P}}_{t,(p_{v-2e_{j}})}\Big{)}=$ $n_{0}D\left(p_{v}\|p_{v-2e_{j}}\right)+\sum_{i=0}^{t-1}D\left(\left(\tilde{P}({p_{v},\hat{P}_{i},\alpha_{i+1}})\right)^{n_{i+1}}\middle\|\left(\tilde{P}({p_{v-2e_{j}},\hat{P}_{i},\alpha_{i+1}})\right)^{n_{i+1}}\middle|\bar{\mathbf{P}}_{i,(p_{v})}\right)$
@@
$\quad+\sum_{i=1}^{t}n_{i}E_{\bar{\mathbf{P}}_{i-1,(p_{v})}}\left[\sum_{j^{\prime}=1}^{k}\left(\alpha_{i}p_{v}[j^{\prime}]+(1-\alpha_{i})\hat{P}_{i-1}[j^{\prime}]\right)\log\frac{\alpha_{i}p_{v}[j^{\prime}]+(1-\alpha_{i})\hat{P}_{i-1}[j^{\prime}]}{\alpha_{i}p_{v-2e_{j}}[j^{\prime}]+(1-\alpha_{i})\hat{P}_{i-1}[j^{\prime}]}\right]$
@@
$=$ $n_{0}\frac{2\delta}{k}\left(\frac{p_{v}[j]}{p_{v-2e_{j}}[j]}-\frac{p_{v}\left[j+\nicefrac{{k}}{{2}}\right]}{p_{v-2e_{j}}\left[j+\nicefrac{{k}}{{2}}\right]}\right)$
@@
$=$ $n_{0}\frac{8\delta^{2}}{k(1-\delta^{2})}+\sum_{i=1}^{t}n_{i}\left(\frac{2\alpha_{i}\delta}{k}\right)^{2}E_{\bar{\mathbf{P}}_{i-1,(p_{v})}}\Bigg{[}\frac{1}{\frac{\alpha_{i}(1-\delta)}{k}+(1-\alpha_{i})\hat{P}_{i-1}[j]}$
$\hskip 184.9429pt+\frac{1}{\frac{\alpha_{i}(1+\delta)}{k}+(1-\alpha_{i})\hat{P}_{i-1}\left[j+\nicefrac{{k}}{{2}}\right]}\Bigg{]},$ (16)
where $(a)$ is a consequence of $\log x\leq x-1$. Recall that $g_{0}(\nicefrac{{1}}{{4k}})=0$ and $\bar{\mathbf{P}}_{i-1,(p_{v})}\left\{\hat{P}_{i-1}[j^{\prime}]<p_{v}[j^{\prime}]-\nicefrac{{1}}{{4k}}\right\}\leq g_{i}(\nicefrac{{1}}{{4k}})$ for every $j^{\prime}\in[1:k]$. Applying Lemma 7 reduces (16) to
$D\Big{(}\bar{\mathbf{P}}_{t,(p_{v})}\Big{\|}\bar{\mathbf{P}}_{t,(p_{v-2e_{i}})}\Big{)}$
$\leq n_{0}\frac{8\delta^{2}}{k(1-\delta^{2})}+\sum_{i=1}^{t}n_{i}\left(\frac{2\alpha_{i}\delta}{k}\right)^{2}\Bigg{(}\frac{1}{\alpha_{i}\frac{1-\delta}{k}+(1-\alpha_{i})\left(\frac{1+\delta}{k}-\frac{1}{4k}\right)}+\frac{\bar{\mathbf{P}}_{i-1,(p_{v})}\left\{\hat{P}_{i-1}[j]<\frac{1+\delta}{k}-\frac{1}{4k}\right\}}{\alpha_{i}\frac{1-\delta}{k}}$
$\hskip 156.49014pt+\frac{1}{\alpha_{i}\frac{1+\delta}{k}+(1-\alpha_{i})\left(\frac{1-\delta}{k}-\frac{1}{4k}\right)}+\frac{\bar{\mathbf{P}}_{i-1,(p_{v})}\left\{\hat{P}_{i-1}\left[j+\frac{k}{2}\right]<\frac{1-\delta}{k}-\frac{1}{4k}\right\}}{\alpha_{i}\frac{1+\delta}{k}}\Bigg{)}$
$\leq n_{0}\frac{8\delta^{2}}{k(1-\delta^{2})}+\sum_{i=1}^{t}n_{i}\frac{(2\alpha_{i}\delta)^{2}}{k}\Bigg{(}\frac{1}{\alpha_{i}({1-\delta})+(1-\alpha_{i})\left(\frac{3}{4}+\delta\right)}+\frac{g_{i}(\nicefrac{{1}}{{4k}})}{\alpha_{i}({1-\delta})}+\frac{1}{\alpha_{i}(1+\delta)+(1-\alpha_{i})\left(\frac{3}{4}-\delta\right)}+\frac{g_{i}(\nicefrac{{1}}{{4k}})}{\alpha_{i}(1+\delta)}\Bigg{)}.$
Now, note that
@@
ensures that $D\left(\bar{\mathbf{P}}_{t,(p_{v})}\middle\|\bar{\mathbf{P}}_{t,(p_{v-2e_{j}})}\right)\leq\nicefrac{{1}}{{2}}$. Using Lemma 6 and substituting (18) and (17) into (14) and (15) concludes the proof. ∎
@@
Once again using $\log x\leq x-1$ gives an expression analogous to (16); the rest of the proof follows the same steps as the proof of Theorem 1. ∎
@@ B-A The order optimal estimator
Combining Lemmas 8 and 9, we arrive at the following intermediate result:
@@
Now, $\alpha-\alpha^{2}\leq\frac{1}{\eta n_{1}}$ ensures that the RHS of (22) is a convex combination of two distributions in $\Delta_{k}$, thus ensuring that $\hat{P}^{(1)}$ is a valid distribution almost surely. Using Lemmas 8 and 9 leads to the desired upper bound on the componentwise variances. ∎
@@ B-B The unprocessed empirical estimator sequence
The proof proceeds inductively. For $t=0$, we have $E[\ell_{2}^{2}(\hat{P}_{0},p)]=\nicefrac{{1}}{{n_{0}}}$. Assume that the estimates at stages $i,0\leq i\leq t-1$ are the empirical distributions $\hat{p}_{\mathrm{emp}}(X_{0}^{n_{0}},\dots,X_{i}^{n_{i}})$ of the samples $X_{i}^{n_{i}}$ respectively, and that the error at stage $t-1$ is bounded above by $\left(\sum_{j}p[j](1-p[j])\right)\left(\sum_{i=0}^{t-1}\frac{n_{i}}{\left(\sum_{j=0}^{i}n_{j}\right)^{2}}\right)$. It is easy to inductively see that each estimator in this sequence is unbiased. When the estimator at stage $t$ is also the empirical estimator of the samples seen up to batch $t$, from (31) and (32), we have
$E[\ell_{2}^{2}(\hat{P}_{t}^{\mathrm{emp}},p)]=$ $\sum_{j\in[1:k]}E\left[\left(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\hat{P}^{\mathrm{emp}}_{t-1}[j]+\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})[j]-p[j]\right)^{2}\right]$
$=$ $\sum_{j\in[1:k]}\Biggl{(}\left(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)+\left(\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})\right)$
$\qquad\qquad+\frac{n_{t}\cdot\sum_{i=0}^{t-1}n_{i}}{\left(\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}\right)^{2}}\cdot\text{cov}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j],\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})\right)\Biggr{)}$
$=$ $\sum_{j\in[1:k]}\Bigg{(}\left(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)$
$\qquad\qquad+\left(\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\left(\frac{p[j](1-p[j])}{n_{t}}+\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)(1-\alpha)^{2}\left(1-\frac{1}{n_{t}}\right)\right)$
$\qquad\qquad+\frac{n_{t}\cdot\sum_{i=0}^{t-1}n_{i}}{\left(\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}\right)^{2}}\cdot\left(2(1-\alpha)\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)\right)\Bigg{)}$
$=$ $\sum_{j\in[1:k]}\left(\text{var}(\hat{P}_{t-1}^{\mathrm{emp}}[j])\left(\left(1-\frac{\alpha_{t}n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}-\frac{(1-\alpha_{t})^{2}n_{t}}{\left(\sum_{i=0}^{t}n_{i}\right)^{2}}\right)+\frac{p[j](1-p[j])n_{t}}{\left(\sum_{i=0}^{t}n_{i}\right)^{2}}\right)$
@@
$\leq$ $\left(\sum_{j}p[j](1-p[j])\right)\cdot\left(\sum_{i=0}^{t}\frac{n_{i}}{\left(\sum_{i^{\prime}=0}^{i}n_{i^{\prime}}\right)^{2}}\right).$
Taking the supremum of $\sum_{j}p[j](1-p[j])$ over $p\in\Delta_{k}$ concludes the proof. ∎

### v4

@@ IV-C1 An example with matching bounds
$=$ $\sum_{i=1}^{t}n_{i}\alpha_{i}\sup_{p}\max_{j}\bar{\mathbf{P}}_{i-1,({p})}\left\{\hat{P}_{i-1}[j]<p[j]{-}\frac{1}{4k}\right\}$
$\leq$ $\sum_{i=1}^{t}a(i+1)^{\beta}\cdot\frac{\text{var}\left(\hat{P}_{i-1}\right)}{\left(\nicefrac{{1}}{{4k}}\right)^{2}}$
$\leq$ $\sum_{i=1}^{t}a(i+1)^{\beta}\cdot 16k^{2}\cdot r_{i-1,k}^{\ell_{2}^{2}}$
$\leq$ $\sum_{i=1}^{t}a(i+1)^{\beta}\cdot 16k^{2}\cdot\frac{\left(1-\frac{1}{k}\right)(\beta{-}\gamma{+}1)}{a\cdot i^{\beta{-}\gamma{+}1}}$
$\leq$ $2^{\beta}\cdot 16k^{2}\cdot(\beta{-}\gamma{+}1)\sum_{i=1}^{t}i^{\gamma-1}$ (8)
@@
This inductively shows that the upper and lower bounds match in terms of order w.r.t. the stage $t$: for each stage $t$, the upper bound is within a constant factor of the lower bound, provided that order-optimal estimators are chosen for stages 0 through t1t{-}1. Moreover, there is a polynomially growing gap between the base and oracle-assisted minimax losses:
@@ IV-C2 An example where the upper and lower bound do not match
We once again note that stronger upper bounds, for example, of the form $h_{t}(\nicefrac{{1}}{{4k}})\leq\exp(-\lambda_{0}t)$, can be assumed by restricting the class of estimators appropriately; this would then make the estimators in Theorem 2 order optimal for every choice of parameters that satisfies the hypothesis condition. Bounds using the Chebyshev inequality remain the tightest without such assumptions.
@@ IV-D General regimes for data accumulation
completing the first part of the proof. If $\nicefrac{{1}}{{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}}=o(\alpha_{t})$, $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$ also diverges (since $\alpha_{t}\leq 1$), and one can find $t_{1}$ such that $g_{t}(\nicefrac{{1}}{{4k}})$ is smaller than $\epsilon$—independent of kk—for $t\geq t_{1}$. The denominator of (12) can then be split into $(1+16k^{2}m)\sum_{i=0}^{t_{1}-1}n_{i}\alpha_{i}^{2}$ and $(1+\epsilon)\sum_{i=t_{1}}^{t}n_{i}\alpha_{i}^{2}$, the latter of which will dominate the former as $t$ grows large. ∎
@@
for each $i$ in [0:t1][0:t-1], for some $M>0$. Then there exists a constant $m>0$ such that the minimax loss at stage $t$ is bounded below as
@@
when the aforementioned estimators are selected for stages 0 to t1t-1.
The proof of this Proposition is deferred to Appendix D. The third condition here is satisfied by all sequences for which $n_{t}\alpha_{t}$ grows exponentially or slower (as can be seen by the Discrete Gronwall inequality [24]). This proposition solidifies the intuition for the lower bound when the mixing probabilities $\alpha_{t}$ decay too quickly: the effective sample size is proportional to the square root of the sample size when in the oracle-assisted setting.
@@ IV-E Data replacement
In this section, we demonstrate the application of Corollary 4 in the data replacement setting where, at every stage $t$, the estimator only has access to samples generated in the current stage $X_{t}^{n}$. Note that $\alpha$ and $n$ do not change with $t$ in this setting, except at stage 0 where all $n$ samples are real. This setting has been explored in depth in the literature e.g. [6, 9, 7, 5, 11, 10, 14, 8].
@@
When $k$ is fixed, we find that since $M\leq\sqrt{2}$ trivially, $m\geq\nicefrac{{c}}{{n\alpha}}$ for some $c\geq 0$. This is especially stark in contrast to the oracle-assisted infimum, which, with probability $1$, is proportional to $\nicefrac{{1}}{{n}}$, corresponding to the case where all samples are real. Thus, in the data replacement setting, the ‘best-case’ minimax expected loss is of the same order as the ‘average’ oracle-assisted loss. This result can also be interpreted as an ‘uncertainty principle’ of recursive estimation in the replacement workflow: the above inequality directly leads to a lower bound on $m\cdot M$.
V Conclusion and Future Work
In this work, we propose and study the minimax discrete distribution estimation problem in self-consuming loops, inspired by recent interest in the empirical machine learning literature. We find minimax bounds on the expected loss, and show the behavior of these bounds in various regimes. We discuss the implications of these results on model collapse considerations, particularly in observing changes in scaling laws.
It should, however, be noted that these results do not directly apply to practical settings. This is in part because we require in our lower bounds that the effective number of samples is comparable to the alphabet size of $X$, which might be very large in practical settings. The class of distributions under consideration in these settings might also be much smaller than the class of distributions considered in the minimax approach. We conjecture that the behavior of the minimax rates in the restricted setting (referred to in the distribution estimation literature as ‘local’ minimax rates) on the number of samples $n_{t}$ and the mixing factors $\alpha_{t}$ remains the same, and the restriction on the space simply changes the dependence on the dimension.
@@ Appendix A Proofs of the minimax lower bounds
We assume that the alphabet size $k$ is even. The worst-case loss for odd $k$ is lower bounded by the worst-case loss for k1k-1, since the supremum is over a smaller set of distributions.
The lower bound argument uses Assouad’s method with the following construction, widely known in the distribution estimation literature e.g. as described in [25, 23, 26]. Consider the set of vectors $\mathcal{V}=\{+1,-1\}^{k/2}$. With each vector $v\in\mathcal{V}$, associate the distribution
@@
The following elementary Lemma will be useful in the proof of Theorem 1:
@@
Using the chain rule for KL divergences, we find the upper bound (16), whenever $j\leq\nicefrac{{k}}{{2}},v[j]=+1$, as follows:
$D\Big(\bar{\mathbf{P}}_{t,(p_{v})}\Big\|\bar{\mathbf{P}}_{t,(p_{v-2e_{j}})}\Big)=$ $n_{0}D\left(p_{v}\|p_{v-2e_{j}}\right)+\sum_{i=1}^{t}D\left(\left(\tilde{P}({p_{v},\hat{P}_{i},\alpha_{i}})\right)^{n_{i}}\middle\|\left(\tilde{P}({p_{v-2e_{j}},\hat{P}_{i},\alpha_{i}})\right)^{n_{i}}\middle|\bar{\mathbf{P}}_{i-1,(p_{v})}\right)$
@@
$+\sum_{i=1}^{t}n_{i}E_{\bar{\mathbf{P}}_{i-1,(p_{v})}}\left[\sum_{j^{\prime}=1}^{k}\left(\alpha_{i}p_{v}[j^{\prime}]+(1-\alpha_{i})\hat{P}_{i-1}[j^{\prime}]\right)\log\frac{\alpha_{i}p_{v}[j^{\prime}]+(1-\alpha_{i})\hat{P}_{i-1}[j^{\prime}]}{\alpha_{i}p_{v-2e_{j}}[j^{\prime}]+(1-\alpha_{i})\hat{P}_{i-1}[j^{\prime}]}\right]$
@@
$\overset{(b)}{=}$ $n_{0}\frac{2\delta}{k}\left(\frac{p_{v}[j]}{p_{v-2e_{j}}[j]}-\frac{p_{v}\left[j+\nicefrac{{k}}{{2}}\right]}{p_{v-2e_{j}}\left[j+\nicefrac{{k}}{{2}}\right]}\right)$
@@
$\overset{(c)}{=}$ $n_{0}\frac{8\delta^{2}}{k(1-\delta^{2})}+\sum_{i=1}^{t}n_{i}\left(\frac{2\alpha_{i}\delta}{k}\right)^{2}E_{\bar{\mathbf{P}}_{i-1,(p_{v})}}\Bigg[\frac{1}{\frac{\alpha_{i}(1-\delta)}{k}+(1-\alpha_{i})\hat{P}_{i-1}[j]}$
$\hskip 184.9429pt+\frac{1}{\frac{\alpha_{i}(1+\delta)}{k}+(1-\alpha_{i})\hat{P}_{i-1}\left[j+\nicefrac{{k}}{{2}}\right]}\Bigg],$
$\overset{(d)}{\leq}$ $n_{0}\frac{8\delta^{2}}{k(1-\delta^{2})}+\sum_{i=1}^{t}n_{i}\left(\frac{2\alpha_{i}\delta}{k}\right)^{2}\Bigg(\frac{1}{\alpha_{i}\frac{1-\delta}{k}+(1-\alpha_{i})\left(\frac{1+\delta}{k}-\frac{1}{4k}\right)}+\frac{1}{\alpha_{i}\frac{1+\delta}{k}+(1-\alpha_{i})\left(\frac{1-\delta}{k}-\frac{1}{4k}\right)}$
$\hskip 142.26378pt+\frac{\bar{\mathbf{P}}_{i-1,(p_{v})}\left\{\hat{P}_{i-1}[j]<\frac{1+\delta}{k}-\frac{1}{4k}\right\}}{\alpha_{i}\frac{1-\delta}{k}}$
$\hskip 142.26378pt+\frac{\bar{\mathbf{P}}_{i-1,(p_{v})}\left\{\hat{P}_{i-1}\left[j+\frac{k}{2}\right]<\frac{1-\delta}{k}-\frac{1}{4k}\right\}}{\alpha_{i}\frac{1+\delta}{k}}\Bigg)$
$\overset{(e)}{\leq}$ $n_{0}\frac{8\delta^{2}}{k(1-\delta^{2})}+\sum_{i=1}^{t}n_{i}\frac{(2\alpha_{i}\delta)^{2}}{k}\Bigg(\frac{1}{\alpha_{i}({1-\delta})+(1-\alpha_{i})\left(\frac{3}{4}+\delta\right)}+\frac{g_{i}(\nicefrac{{1}}{{4k}})}{\alpha_{i}({1-\delta})}$
$\hskip 142.26378pt+\frac{1}{\alpha_{i}(1+\delta)+(1-\alpha_{i})\left(\frac{3}{4}-\delta\right)}+\frac{g_{i}(\nicefrac{{1}}{{4k}})}{\alpha_{i}(1+\delta)}\Bigg).$ (16)
Here, $(a)$ is a consequence of $\log x\leq x-1$, $(b)$ and $(c)$ follow from the construction of $p_{v}$, $(d)$ follows from Lemma 8, and $(e)$ follows from Definition 3. Now, note that
@@
ensures that $D\left(\bar{\mathbf{P}}_{t,(p_{v})}\middle\|\bar{\mathbf{P}}_{t,(p_{v-2e_{j}})}\right)\leq\nicefrac{{1}}{{2}}$. Invoking Lemma 7 and substituting (18) and (17) into (14) and (15) concludes the proof. ∎
@@
The rest of the proof follows the same steps as the proof of Theorem 1. ∎
@@ B-A The order optimal estimator
Combining Lemmas 9 and 10, we arrive at the following intermediate result:
@@
Now, $\alpha-\alpha^{2}\leq\frac{1}{\eta n_{1}}$ ensures that the RHS of (22) is a convex combination of two distributions in $\Delta_{k}$, thus ensuring that $\hat{P}^{(1)}$ is a valid distribution almost surely. Using Lemmas 9 and 10 leads to the desired upper bound on the componentwise variances. ∎
@@ B-B The unprocessed empirical estimator sequence
The proof proceeds inductively. For $t=0$, we have $E[\ell_{2}^{2}(\hat{P}_{0},p)]=\nicefrac{{1}}{{n_{0}}}$. Assume that the estimates at stages $i,0\leq i\leq t-1$ are the empirical distributions $\hat{p}_{\mathrm{emp}}(X_{0}^{n_{0}},\dots,X_{i}^{n_{i}})$ of the samples $X_{i}^{n_{i}}$ respectively, and that the error at stage t1t-1 is bounded above by $\left(\sum_{j}p[j](1-p[j])\right)\left(\sum_{i=0}^{t-1}\frac{n_{i}}{\left(\sum_{j=0}^{i}n_{j}\right)^{2}}\right)$. It is easy to inductively see that each estimator in this sequence is unbiased by observing that the expected value of the empirical distribution at stage $t$ equals $\alpha_{t}p+(1-\alpha_{t})\hat{P}_{t-1}$.
When the estimator at stage $t$ is also the empirical estimator of the samples seen up to batch $t$, we have the upper bound (23) as follows:
$E[\ell_{2}^{2}(\hat{P}_{t}^{\mathrm{emp}},p)]=$ $\sum_{j\in[1:k]}E\left[\Bigg(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\hat{P}^{\mathrm{emp}}_{t-1}[j]+\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})[j]-p[j]\Bigg)^{2}\right]$
$=$ $\sum_{j\in[1:k]}\Biggl(\left(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)+\left(\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})[j]\right)$
$\qquad\quad+\frac{n_{t}\cdot\sum_{i=0}^{t-1}n_{i}}{\left(\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}\right)^{2}}\cdot\text{cov}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j],\hat{p}_{\mathrm{emp}}(X_{t}^{n_{t}})[j]\right)\Biggr)$
$\overset{(a)}{=}$ $\sum_{j\in[1:k]}\Bigg(\left(\frac{\sum_{i=0}^{t-1}n_{i}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)+\frac{n_{t}\cdot\sum_{i=0}^{t-1}n_{i}}{\left(\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}\right)^{2}}\cdot\left(2(1-\alpha)\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)\right)$
$\qquad\quad+\left(\frac{n_{t}}{\sum_{i^{\prime}=0}^{t}n_{i^{\prime}}}\right)^{2}\cdot\Bigg(\frac{p[j](1-p[j])}{n_{t}}+\text{var}\left(\hat{P}^{\mathrm{emp}}_{t-1}[j]\right)(1-\alpha)^{2}\left(1-\frac{1}{n_{t}}\right)\Bigg)\Bigg)$
$=$ $\sum_{j\in[1:k]}\Bigg(\text{var}(\hat{P}_{t-1}^{\mathrm{emp}}[j])\cdot\left(\left(1{-}\frac{\alpha_{t}n_{t}}{\sum_{s{=}0}^{t}n_{s}}\right)^{2}{-}\frac{(1{-}\alpha_{t})^{2}n_{t}}{\left(\sum_{i{=}0}^{t}n_{i}\right)^{2}}\right)+\frac{p[j](1-p[j])n_{t}}{\left(\sum_{i=0}^{t}n_{i}\right)^{2}}\Bigg)$
@@
$\leq$ $\left(\sum_{j}p[j](1{-}p[j])\right)\cdot\left(\sum_{i=0}^{t}\frac{n_{i}}{\left(\sum_{i^{\prime}=0}^{i}n_{i^{\prime}}\right)^{2}}\right).$ (23)
The expressions for the variance of $\hat{p}^{\mathrm{emp}}(.)$ and its covariance with $\hat{P}^{\mathrm{emp}}_{t-1}$ used for step $(a)$ are derived in (32) and (33) respectively. Taking the supremum of $\sum_{j}p[j](1-p[j])$ over $p\in\Delta_{k}$ concludes the proof. ∎